# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v3

@@ 4 Data description
The effectiveness of real-time COR estimation is critically dependent on the fidelity and coverage of trajectory data. Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020). However, these systems suffer from narrow fields of view, occlusion, overlapping detections, and high maintenance costs, which limit scalability and data quality (St-Aubin et al., 2013). Recent advances in AV sensing offer a promising alternative. AV datasets provide high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers. This richness allows for realistic near-miss detection in complex traffic environments. Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation. For example, studies using the Waymo Open Dataset have focused mainly on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios that span 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions, such as intersection near-misses, lane changes, and pedestrian crossings, conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite comprises seven ring cameras and two front-facing stereo cameras (20 Hz), supported by dual 64-beam VLP-32C LiDAR sensors, which produce  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
@@
Argoverse-2 was selected after benchmarking open-source AV datasets (Table 1). Waymo Open offers a greater total mileage but lower near-miss density and no AV-behavior labels. Lyft Level-5 is confined to a single city and lacks scene-level filtering. nuScenes samples at 2 Hz, limiting the capture of rapid interactions. In contrast, Argoverse-2 balances spatial coverage, sensor fidelity, near-miss-rich scenarios, and behavioral annotations, making it well-suited to this study’s goals. This study used motion-forecasting and HD-map modules of Argoverse-2 to extract vehicle trajectories, compute near-miss indicators, and encode roadway geometry.
4.1 Data Preprocessing
To operationalize the proposed framework, both the motion forecasting and HD map modules of the Argoverse-2 dataset were processed through a structured pipeline. The preprocessing consisted of four key stages: (i) extraction of multi-agent trajectories, (ii) transformation of coordinates from local map space to global geographic space, (iii) integration with HD map features to contextualize V-I interactions, and (iv) selection of a high-risk urban corridor for detailed analysis.
4.1.1 Motion forecasting dataset
@@
Each motion forecasting scenario follows a hierarchical structure with three levels: scenario metadata, agent tracks, and frame-level states. Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling. At the scenario level, metadata includes a unique scenario ID, the recording city, and the focal agent, typically the ego AV. Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states. Frame-level states capture position, velocity, and heading at each timestep, forming the basis for computing 2D-TTC.
@@
To implement the framework, we developed a Python-based batch pipeline using the Argoverse-2 SDK. The pipeline parsed motion forecasting scenarios across all cities and extracted tracks for all agents—not just the focal vehicle—to retain critical multi-agent interactions. For each track, frame-level kinematic attributes were retrieved, including $(x,y)$ position, heading, velocity components, and timestamps.
Because positions were reported in local, city-specific frames, all coordinates were transformed into a globally referenced WGS84 system. This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj. This geospatial normalization provided consistent referencing across cities and ensured compatibility with HD map overlays. Batch scripts automatically traversed scenario folders, performed attribute extraction and coordinate conversion, and exported unified CSV datasets containing both original city-frame and converted lat/lon values. These datasets were then used for spatial filtering, interaction detection, and overlay validation.
The processed files formed the basis for V–V and V–I near-miss detection. For each valid scenario, pairwise distances and relative approach rates were computed for all vehicle pairs, while V-I distances were derived from buffered polygons of lane edges and curbs in the HD map. Scenarios with no substantive interaction (e.g., straight-lane cruising without lane changes or turns) were excluded, ensuring the dataset emphasized behaviorally informative, high-risk conditions consistent with proactive safety modeling. Non-vehicle agents (e.g., pedestrians, cyclists) were also excluded at this stage, though the pipeline is extensible to incorporate vulnerable road users in future work.
4.1.2 HD map
@@
Because all coordinates are expressed in city-specific Cartesian frames, global georeferencing was required to align HD maps with trajectory data. For each city, the WGS84 geodetic origin was projected into its respective UTM zone (e.g., Zone 17 for Miami and Pittsburgh, Zone 14 for Austin, Zone 10 for Palo Alto, Zone 18 for Washington, D.C.) using the pyproj library. Each local $(x,y)$ coordinate was then offset by this origin and inverse-projected into latitude–longitude, ensuring spatial consistency across cities. Once transformed, all HD map components: lane boundaries, centerlines, and crosswalks—were indexed using their global coordinates and converted into shapefiles or GeoPandas geometries. This geospatial transformation enabled efficient spatial querying, collision proximity analysis, and visual overlay with satellite imagery and trajectory data. In particular, the lane boundaries were instrumental in detecting V–I near-misses by allowing for the computation of lateral vehicle encroachments beyond the legal driving space. From the processed dataset, a high-density urban corridor in Miami was present to showcase the transformation (see Fig. 4.1.2). This illustrates the transformation of AV data into real-world coordinates, its alignment with HD maps and satellite imagery, and the resulting traffic patterns that highlight interaction hotspots across intersections and midblocks.
Fig. 3.: AV data transformed to real-world coordinates and aligned: (a) trajectories with lane boundaries and crosswalks, (b) validation with satellite imagery, and (c–f) directional traffic patterns highlighting interaction hotspots.
4.1.3 Trajectory preprocessing and derivation of vehicle dynamics
Before computing 2D-TTC, a series of preprocessing steps was applied to ensure the precision, continuity, and physical plausibility of vehicle motion data from the Argoverse-2 dataset. These steps included the extraction of raw trajectory data, trajectory smoothing and filtering, and the derivation of dynamic variables such as speed, acceleration, and steering angle. First, all multi-agent trajectories were transformed from local city-specific frames to global WGS84 coordinates, enabling consistent alignment with HD map geometries. Raw trajectories were then examined for noise caused by sensor jitter, occlusion, or missed detections—issues more common in human-driven vehicles than in AV tracks. To suppress such noise while preserving curvature and lane-change dynamics, cubic B-spline smoothing was applied to the position coordinates (Eilers and Marx, 1996; Choi et al., 2024). This method effectively removed local fluctuations while retaining key geometric features (see Fig.4.1.3). Stationary or irrelevant agents, such as parked vehicles and bicycles, were filtered out to reduce computational load and focus only on interaction-relevant entities.
@@ 4.2 Study area
Under realistic and heterogeneous urban traffic conditions, a representative high-volume corridor was selected along Biscayne Boulevard in Miami, Florida. This arterial extends from Downtown Miami through the Upper East Side toward the northern city limits, serving as a critical connector to major activity centers, including the Miami International Airport region. It is characterized by multimodal interactions, geometric complexity, frequent turning maneuvers, and diverse traffic control features. The corridor was chosen for its rich coverage in the Argoverse-2 motion forecasting dataset, HD map, and its suitability for extracting sufficient samples for extreme near-miss scenarios.
As shown in Fig. 4.2, the study area was extracted using ArcGIS Pro based on scenario density, HD map fidelity, and heterogeneity of vehicle maneuvers. The extracted network includes a continuous corridor with multiple intersections, midblock segments, and bidirectional traffic operations. For analytical consistency, the corridor was subdivided into nine directional segments (S1–S9) and ten intersections (I1–I10). Segment boundaries were defined midblock to maintain flow continuity, while intersections were delineated to capture near-miss-prone zones such as turning paths, crosswalks, and lane merges. Each subregion was mapped with lane boundaries, crosswalks, and vehicle trajectories to preserve geometric fidelity and directional alignment. Insets in Fig. 4.2 illustrate the spatial configuration of the segmented units, which serve as the basis for grouped modeling and localized risk estimation.
@@
Table 2 provides a detailed profile of the 28 study subregions, consisting of 10 intersections and 18 directional segments. Intersections generally show fewer scenarios and unique vehicles (e.g., I2 with 47 scenarios and 220 vehicles, I4 with 96 scenarios and 568 vehicles) compared to adjacent midblock segments, reflecting shorter dwell times but higher near-miss intensity. In contrast, segments accumulate a significantly higher number of scenarios and vehicles due to sustained travel. For example, S7 (NE 19-17 St) records 308 northbound and 332 southbound scenarios involving more than 4000 vehicles, while S8 (NE 17–15 St) captures nearly 4400 vehicles in 608 scenarios.
@@
Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design. Driveway and minor intersection density also differ across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses. The median type alternates between raised (e.g., S1–S2) and flush-paved (e.g., S3–S9), influencing turning maneuvers and mid-block encroachments. Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
Building on the spatial segmentation and detailed subregion profiles outlined above, the next step involves quantifying dynamic safety conditions using the proposed 2D-TTC framework (Algorithm 1). For each block, extreme near-miss events were identified by forward simulating vehicle trajectories under the geometric and traffic conditions described in Table 2. Both V–V and V–I interactions were modeled to reflect the heterogeneous near-miss environments of intersections and segments. Trajectories were projected over a 3s horizon, and spatial overlaps with surrounding traffic or roadway boundaries were detected as potential near-miss events. Representative outcomes from this process are shown in Fig. 4.2, illustrating the spatial dynamics of predicted V–V and V–I near-misses within the Miami study corridor. These event visualizations provide the empirical basis for block-level 2D-TTC extraction across the 10 intersections (I1–I10), enabling localized evaluation of crash-prone conditions.
Table 3: Summary statistics for V–V and V–I near-misses at Intersection
@@
TTC 1.617 0.11 2.99 0.863 1.304 0.10 2.99 0.889
Relative speed 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
Relative acc/dec -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
Relative distance 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
Jerk -5.550 -13.609 9.315 2.652 -7.900 -10.800 2.800 14.700
Heading difference 1.077 0.00 6.251 1.410 -0.045 -3.138 3.141 1.647
Steering difference 0.409 0.00 3.142 0.666 -0.006 -1.571 1.571 0.130
Volume 13.462 2.00 31.00 5.969 10.461 1.00 31.00 5.392
Vehicle length 4.78 3.18 6.92 0.49 5.1 3.95 12.65 1.30
Vehicle width 2.22 1.76 3.16 0.30 2.04 1.36 3.08 0.23
@@
Descriptive statistics of intersection-level V–V and V–I near-misses are summarized in Table 3. The continuous indicators include 2D-TTC, relative speed, acceleration, distance, jerk, and angular deviations (heading and steering), along with vehicle dimensions and exposure (volume). On average, 2D-TTC values for both near-miss types range between 1.3 and 1.6 s, with extreme cases falling below 0.1 s, reflecting highly critical near-miss conditions. V–I near-misses are associated with higher relative speeds and tighter distance gaps, indicating more abrupt encroachment against roadway boundaries. Conversely, V–V near-misses exhibit larger heading and steering deviations, along with higher jerk variability, suggesting unstable maneuvers such as sudden turns or evasive swerves. The categorical distribution of near-misses further illustrates key behavioral differences. Nearly three-quarters (72%) of V–V events occurred during through movements, with fewer during left (26%) or right (9%) turns. In contrast, V–I near-misses were almost entirely turning-related, with 46% occurring in left turns and 54% in right turns, and none during through movements. This pattern highlights that infrastructure-related near-misses typically occur during lateral maneuvers at intersections, where vehicles cross lane boundaries or encroach on curbs and medians.
Beyond intersections, the analysis was extended to directional roadway segments (S1–S9) using the same 2D-TTC framework (See Fig. 4.2).Table 4 summarizes descriptive statistics of both V–V and V–I near misses at the segment level. For V–V near-misses, average TTC values were longer (1.83 s) than those observed at intersections, suggesting greater temporal headway between vehicles in segment settings. However, the extreme minima (0.11 s) and high maximum relative speeds (up to 33.9 m/s) highlight the presence of aggressive overtaking and high-speed convergence. Relative distances also exhibited a much broader range (up to 141.5 m), indicating that near-misses often arise in more dispersed longitudinal interactions rather than tightly constrained intersection maneuvers.
@@
TTC 1.829 0.11 2.99 0.805 1.272 0.10 2.99 0.912
Relative speed 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
Relative acc/dec -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
Relative distance 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02
Jerk -7.121 -17.502 8.423 1.552 -5.70 -12.72 3.556 4.77
Heading difference 0.368 0.000 6.256 0.990 -0.032 -3.102 3.138 1.586
Steering difference 0.577 0.000 3.142 0.756 -0.001 -1.571 1.571 0.167
Volume 16.801 1.000 65.000 11.009 12.816 1.000 65.000 7.906
Vehicle Length 4.74 3.0 7.85 1.11 4.79 2.95 8.65 0.87
Vehicle width 2.12 1.36 3.08 0.23 2.22 1.76 4 0.18
@@
For V–I near-misses, the mean TTC was shorter (1.27s), with relative speeds peaking at 28.1 m/s, underscoring the abrupt nature of vehicle–boundary encounters. Compared with intersections, segment-based V–I near-misses involved longer approach distances but also showed greater instability in jerk and heading, reflecting frequent lateral drifts toward lane edges or curbs. These events often occurred near driveways and minor access points, where vehicles deviated from their intended path, creating encroachment risks. The categorical breakdown reinforces these patterns. For V–V events, nearly 59% of near-misses were associated with through movements, while about 32% and 10% were linked to right and left turns, respectively. In contrast, V–I near-misses were dominated by turning maneuvers, with more than half (53.9%) linked to right turns and 46% to left turns, and no near-misses detected during through movements. This distribution highlights the heightened role of lateral maneuvers in producing infrastructure-related risks. Furthermore, median type also influenced outcomes: 76% of V–V near-misses and 86% of V–I near-misses occurred along undivided sections, underscoring the vulnerability of open medians and flush-paved designs in segment environments.
Taken together, these segment-level findings reveal that while intersections are hotspots for tightly constrained V-V near-misses, segments carry distinctive risks tied to high-speed encounters and infrastructure encroachments. The observed distributions underscore the importance of treating intersections and segments as functionally distinct safety domains when modeling near-miss dynamics and designing targeted countermeasures.

### v4

@@ 4 Data description
The effectiveness of real-time COR estimation depends on the spatiotemporal fidelity and contextual coverage of trajectory data. Traditional sensing technologies (e.g., roadside cameras, loop detectors, fixed LiDAR) can support safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020). However, these systems often face limited spatial coverage, occlusion, overlapping detections, and deployment/maintenance burdens, which constrain scalability and data quality (St-Aubin et al., 2013). In contrast, AV datasets provide high-frequency, multi-agent trajectories with precise localization, enabling consistent reconstruction of short-horizon kinematics and interaction geometry needed for near-miss-based COR modeling. Despite this potential, AV data remain underutilized in generalized real-time COR frameworks. For example, studies using the Waymo Open Dataset have mainly analyzed localized conflicts and often omit HD-map-aligned roadway boundaries needed to represent V–I exposure, limiting transferability to corridor-scale inference (Anis et al., 2025b).
To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and Washington, D.C.) using SAE Level 4 AVs. Argoverse-2 contains >>250k driving scenarios spanning 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is curated to emphasize interaction-rich episodes (e.g., intersection encounters, lane changes, and complex crossings) that are informative for extreme-conflict modeling. A defining feature is its scene-adaptive sampling, which prioritizes high-interaction moments rather than routine free-flow conditions. Because this design intentionally over-samples kinematically and socially unusual behavior (Wilson et al., 2023), the extracted near-misses and inferred COR should be interpreted as conditional on safety-critical traffic states, rather than as unconditional averages over typical corridor operations. Scenarios include agent identifiers, AV status, and map-based behavioral annotations. The sensor suite comprises seven ring cameras and two front-facing stereo cameras (20 Hz), supported by dual 64-beam VLP-32C LiDAR sensors, which produce  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas, all registered to local coordinates.
@@
Argoverse-2 was selected after benchmarking open-source AV datasets (Table 1). Waymo Open offers greater overall mileage but does not explicitly target interaction-dense segments at the same level, Lyft Level-5 is limited to a single city, and nuScenes samples at 2 Hz, reducing sensitivity to rapid interactions. Argoverse-2 offers a practical balance of sampling rate, multi-agent density, HD-map detail, and interaction-centric scenario selection, making it well-suited for geometry-aware near-miss extraction and corridor-scale COR estimation. This study used vehicle trajectories and HD-map details, which provide multiple agent tracks labeled by object type (e.g., vehicle, pedestrian, cyclist). To operationalize the proposed framework, we processed these modules through a four-stage pipeline: (i) extraction of multi-agent trajectories and necessary information, (ii) coordinate transformation from local map space to global geographic space, (iii) integration with HD-map boundary features to define V–I exposure, and (iv) selection of a high-risk urban corridor for detailed analysis.
4.1 Motion forecasting dataset
@@
Each Argoverse-2 motion-forecasting scenario has a three-level structure: scenario metadata, agent tracks, and frame-level states. Scenarios are 11 s long and provide a birds-eye-view 2D centroid and heading for each tracked object at 10 Hz (110 time steps), enabling fine-grained interaction modeling. Scenario metadata includes a unique scenario ID, recording city, and a focal agent (typically the ego AV). Each scenario contains multiple agent tracks with object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal/scored/unscored), and a time-ordered sequence of states. Frame-level states provide position, velocity components, and heading, which form the basis for computing 2D-TTC. We directly use these trajectories as provided by Argoverse-2; tracks are produced by Argo’s internal multi-sensor fusion and map-aligned ego poses, which help keep trajectories tightly registered to the HD map.
@@
To implement the proposed framework, we developed a Python batch-processing pipeline using the Argoverse-2 SDK. The pipeline parses scenarios and extracts tracks for all agents (not only the focal vehicle) to preserve the multi-agent interaction context. For each track, we retrieve frame-level kinematic attributes, including $(x,y)$ position, heading, velocity components, and timestamps. For the present analysis, we retain only motor-vehicle agents and filter out pedestrians and cyclists; the same pipeline can be extended to incorporate vulnerable road users in future work.
Because positions are reported in city-specific local coordinate frames, all coordinates are transformed to a globally referenced WGS84 system. This is done by projecting each city’s local origin into the appropriate UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), offsetting local $(x,y)$ values, and inverse-projecting to latitude/longitude using pyproj. Batch scripts traverse scenario folders, extract attributes and convert coordinates, and export unified CSV files containing both the original city-frame coordinates and the converted lat/lon values. To verify that this transformation preserves trajectory geometry, we conducted a visual geospatial validation for the Biscayne Boulevard corridor by overlaying converted trajectories and HD-map features on satellite imagery (Fig. 4.2). Vehicle paths closely follow mapped lane centerlines and intersection geometry, indicating residual registration error is small relative to lane width and is acceptable for conflict localization.
In addition to the motion-forecasting trajectories, we use the Argoverse-2 sensor annotation files $(\texttt{annotations.feather})$, which provide 3D bounding-box dimensions for each tracked object at each timestamp (including length_m and width_m). The motion-forecasting tracks are joined to these annotations using the shared scenario identifier, timestamp_ns, and track_uuid. For each vehicle track, the corresponding length_m and width_m values are used as the plan-view dimensions of the oriented rectangular footprint in the geometry-aware 2D-TTC computation.
The processed files form the input for V–V and V–I near-miss detection. For each valid scenario, we evaluate V-V interactions by computing pairwise relative spacing and approach rates across vehicle pairs, while V–I proximity is evaluated relative to HD-map-derived lane edges/curb boundaries. Scenarios without substantive interaction (e.g., sustained straight-lane cruising without turning, merging, or lateral maneuvers) are excluded to emphasize behaviorally informative, high-interaction conditions consistent with proactive extreme-conflict modeling.
4.2 HD map
@@
Because all coordinates are expressed in city-specific Cartesian frames, global georeferencing was required to align HD maps with trajectory data. For each city, the WGS84 geodetic origin was projected into its respective UTM zone (e.g., Zone 17 for Miami and Pittsburgh, Zone 14 for Austin, Zone 10 for Palo Alto, Zone 18 for Washington, D.C.) using the pyproj library. Each local $(x,y)$ coordinate was then offset by this origin and inverse-projected into latitude–longitude, ensuring spatial consistency across cities. Once transformed, all HD map components: lane boundaries, centerlines, and crosswalks—were indexed by their global coordinates and converted into shapefiles or GeoPandas geometries. This geospatial transformation enabled efficient spatial querying, collision proximity analysis, and visual overlay with satellite imagery and trajectory data. In particular, the lane boundaries were instrumental in detecting V–I near-misses by allowing for the computation of lateral vehicle encroachments beyond the legal driving space. From the processed dataset, a high-density urban corridor in Miami was present to showcase the transformation (see Fig. 4.2). This illustrates the transformation of AV data into real-world coordinates, its alignment with HD maps and satellite imagery, and the resulting traffic patterns that highlight interaction hotspots across intersections and midblocks.
Fig. 3.: AV data transformed to real-world coordinates and aligned: (a) trajectories with lane boundaries and crosswalks, (b) validation with satellite imagery, and (c–f) directional traffic patterns highlighting interaction hotspots
4.3 Trajectory preprocessing and derivation of vehicle dynamics
Before computing 2D-TTC, a series of preprocessing steps was applied to ensure the precision, continuity, and physical plausibility of vehicle motion data from the Argoverse-2 dataset. These steps included extracting raw trajectory data, smoothing and filtering the trajectories, and deriving dynamic variables such as speed, acceleration, and steering angle. First, all multi-agent trajectories were transformed from local city-specific frames to global WGS84 coordinates, enabling consistent alignment with HD map geometries. Raw trajectories were then examined for noise caused by sensor jitter, occlusion, or missed detections—issues more common in human-driven vehicles than in AV tracks. To suppress such noise while preserving curvature and lane-change dynamics, cubic B-spline smoothing was applied to the position coordinates (Eilers and Marx, 1996; Choi et al., 2024). This method effectively removed local fluctuations while retaining key geometric features (see Fig. ref fig:4). Stationary or irrelevant agents, such as parked vehicles and bicycles, were filtered out to reduce computational load and focus only on interaction-relevant entities.
@@ 4.4 Study area
Under realistic and heterogeneous urban traffic conditions, we selected a high-volume study corridor along Biscayne Boulevard in Miami, Florida. This arterial links Downtown Miami to the Upper East Side and northern city limits and is characterized by dense access activity, frequent turning and lane-changing maneuvers, and diverse traffic control. The corridor was chosen because it has high scenario density in Argoverse-2 and high-fidelity HD map coverage, and because its mapped curb and median boundaries provide explicit roadway-edge geometry needed to consistently define V–I near-misses.
As shown in Fig. 4.4, the study area was extracted in ArcGIS Pro based on scenario density, HD map fidelity, and maneuver diversity. The resulting network forms a continuous corridor with bidirectional operations, multiple intersections, and midblock sections. For analytical consistency, the corridor was subdivided into 10 intersections (I1–I10) and 18 directional segments organized into nine segment pairs (S1–S9). Segment limits were defined midblock to preserve operational continuity, while intersection extents were delineated to capture conflict-prone subareas (turning paths, merges, and channelization transitions). These subregions define the hierarchical modeling structure and enable direction-aware inference.
@@
Table 2 summarizes the 28 study subregions. Intersections generally include fewer scenarios and unique vehicles than adjacent segments (e.g., I2: 47 scenarios/220 vehicles; I4: 96 scenarios/568 vehicles), reflecting shorter dwell time within intersection boundaries. By contrast, segments accumulate substantially more scenarios and vehicles due to sustained travel (e.g., S7 and S8 each include several hundred scenarios and >>4000 unique vehicles)
@@
Lane configurations range from 2 to 3 lanes per direction, with average lane widths of 3.1 m (S7) to 3.6 m (S6). Access density also varies across segments (e.g., multiple driveways/minor access points in S3–S4 versus few or none in S5–S6), creating heterogeneous exposure to lateral disturbances and side-entry maneuvers. Median treatments alternate between raised and flush-paved sections, which can affect turning opportunities and the potential for edge encroachment.
Building on the spatial segmentation and subregion profiles above, we next quantify time-varying safety conditions using the proposed 2D-TTC pipeline (Algorithm 1). Because this study adopts the BM sampling strategy, there is no global consensus on the optimal block duration. Prior BM–GEV studies based on video data often employ long, time-based blocks (e.g., several minutes) and select block sizes based on goodness-of-fit criteria. In contrast, AV trajectory datasets consist of short scenarios (e.g., Argoverse-2: 17 s; Waymo: 20 s; KITTI: 1–2 min) and high-frequency interaction windows. Consequently, this study adopts an interaction-window BM strategy, consistent with real-time near-miss modeling in our prior work (Anis et al., 2025b), in which blocks are defined by unique interaction units (Ali et al., 2022a). Specifically, each continuously observed interaction window, either a V–V or a V–I interaction within a scenario, is treated as a block, and the block extreme is defined as the minimum 2D-TTC over the 11 s (110-frame) window. For each interaction window, the near-miss process is evaluated at every frame by forward-simulating motion using an RK4-based kinematic projection and subsequently computing 2D-TTC. Thus, each retained extreme summarizes 110 high-frequency kinematic states and 110 interaction assessments, providing substantial within-block information despite the short scenario duration.
Because multiple vehicles are present in a single 11s scenario, multiple interaction windows occur simultaneously. We therefore distinguish between (i) candidate near-miss detected across vehicles and frames within a scenario and (ii) the single extreme block retained for EVT per interaction window. Defining blocks by interaction windows also avoids repeatedly selecting adjacent frame-wise minima from the same persistent interaction, which would overweight long-lasting conflicts and amplify temporal dependence. For each retained block, extreme near-miss events are identified by forward-simulating vehicle trajectories under the geometric and traffic conditions summarized in Table 2. Trajectories are projected over a 3s horizon, and spatial overlaps with surrounding traffic or roadway boundaries are evaluated as potential near-miss events. To focus the BM sample on genuinely unsafe edge-encroachment behavior, we retain V–I encounters only when the interaction window contains short 2D-TTC values with moderate-to-high approach speeds, filtering out routine low-speed turning proximity that can appear as boundary approach under constant-input projection. Representative outcomes from this process are shown in Fig. 4.4, illustrating the spatial dynamics of predicted V–V and V–I near-misses within the Miami study corridor. These event visualizations provide the empirical basis for block-level 2D-TTC extraction across the 10 intersections (I1–I10), enabling localized evaluation of crash-prone conditions. Finally, all microscopic covariates (e.g., relative speed, relative acceleration/deceleration, jerk, heading/steering differences) are extracted from the same interaction window and aligned to the extreme event, either evaluated at the frame where the extreme 2D-TTC occurs.
Table 3: Summary statistics for V–V and V–I near-misses at Intersections
@@
TTC (s) 1.617 0.11 2.99 0.863 1.304 0.10 2.99 0.889
Relative speed (m/s) 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
Relative acc./dec. (m/s2) -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
Relative distance (m) 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
Jerk ($\mathrm{m/s^{3}}$) -5.550 -13.609 9.315 2.652 -7.900 -10.800 2.800 14.700
Heading difference (rad) 1.077 0.00 6.251 1.410 -0.045 -3.138 3.141 1.647
Steering difference (rad) 0.409 0.00 3.142 0.666 -0.006 -1.571 1.571 0.130
Traffic volume (veh) 13.462 2.00 31.00 5.969 10.461 1.00 31.00 5.392
Vehicle length (m) 4.78 3.18 6.92 0.49 5.1 3.95 12.65 1.30
Vehicle width (m) 2.22 1.76 3.16 0.30 2.04 1.36 3.08 0.23
@@
Descriptive statistics for intersection-level V–V and V–I near-misses are reported in Table 3. The continuous indicators include 2D-TTC and interaction dynamics (relative speed, relative acceleration/deceleration, relative distance, jerk, and heading/steering differences), along with vehicle dimensions and exposure (volume). Mean 2D-TTC values are approximately 1–2s (V–V: 1.62s; V–I: 1.30s), while the most safety-critical cases lie in the lower tail, with minima near 0.10–0.11s. Table further quantifies the lower-tail structure of 2D-TTC, showing that crash-prone near-miss risk is concentrated in a narrow extreme range. To ensure that extremely small 2DTTC values from geometry-oriented rectangle footprints (e.g., spurious corner overlap), the corresponding scenarios were manually reviewed by visualizing the time-evolving oriented bounding boxes and headings; the critical cases reflected plausible relative motion and geometry rather than numerical overlap errors. V–I near-misses tend to exhibit higher relative speeds and smaller distance gaps, consistent with abrupt boundary approaches, whereas V–V near-misses show larger heading/steering deviations and higher jerk variability, consistent with unstable maneuvers (e.g., abrupt turns or evasive swerves). Categorical summaries further highlight differences: most V–V events occur during through movements, whereas V–I near-misses are concentrated in lateral/turning maneuvers, consistent with edge approach under turning and merge dynamics.
The analysis was also conducted for directional segments (S1–S9) using the same 2D-TTC pipeline (See Fig. 4.4). Table 4 reports segment-level descriptive statistics. For V–V near-misses, mean TTC values are longer than at intersections (1.83 s), suggesting larger headways in midblock traffic, while very small TTC minima and high relative-speed maxima indicate occasional aggressive overtaking and rapid convergence. Relative distances span a wider range than at intersections, reflecting more dispersed longitudinal interactions in segment settings.
@@
TTC (s) 1.829 0.11 2.99 0.805 1.272 0.10 2.99 0.912
Relative speed ($\mathrm{m/s}$) 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
Relative acc./dec. (m/s2) -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
Relative distance (m) 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02
Jerk ($\mathrm{m/s^{3}}$) -7.121 -17.502 8.423 1.552 -5.70 -12.72 3.556 4.77
Heading difference (rad) 0.368 0.000 6.256 0.990 -0.032 -3.102 3.138 1.586
Steering difference (rad) 0.577 0.000 3.142 0.756 -0.001 -1.571 1.571 0.167
Traffic volume (veh) 16.801 1.000 65.000 11.009 12.816 1.000 65.000 7.906
Vehicle Length (m) 4.74 3.0 7.85 1.11 4.79 2.95 8.65 0.87
Vehicle width (m) 2.12 1.36 3.08 0.23 2.22 1.76 4 0.18
Driveway intensity ($\mathrm{m^{-1}}$) 0.009 0.000 0.015 0.005 0.009 0.000 0.015 0.005
Lane width (m) 3.28 2.99 3.96 0.14 3.27 2.99 3.96 0.141
@@
For V–I near-misses, the mean TTC was shorter with greater relative speeds, underscoring the abrupt nature of vehicle–infrastructure encounters. Compared with intersections, segment-level V–I near-misses tend to occur at longer approach distances but exhibit greater variability in jerk and heading, consistent with lateral drift or corrective steering as vehicles move toward lane edges, curbs, or medians. These events frequently cluster near driveways and minor access points, where side-entry activity and path adjustments increase the likelihood of encroachment. Because the V–I definition is restricted to same-carriageway curbs/medians and the retained encounters are filtered by 2D-TTC, the resulting V–I sample reflects elevated boundary-encroachment risk rather than generic turning or lane-change exposure. Segment-level access exposure is captured by driveway intensity, defined as the density of driveways and minor access points per unit segment length. The categorical breakdown reinforces these patterns. For V–V events, nearly 59% of near-misses were associated with through movements, while about 32% and 10% were linked to right and left turns, respectively. In contrast, V–I near-misses were dominated by turning maneuvers, with more than half (53.9%) linked to right turns and 46% to left turns, and no near-misses detected during through movements. This distribution highlights the heightened role of lateral maneuvers in producing infrastructure-related risks. Furthermore, median type also influenced outcomes: 76% of V–V near-misses and 86% of V–I near-misses occurred along undivided sections, underscoring the vulnerability of open medians and flush-paved designs in segment environments.
Taken together, these segment-level results indicate that, while intersections concentrate tightly constrained V–V conflicts, segments exhibit distinct risk mechanisms associated with higher-speed interactions and boundary encroachments. This supports treating intersections and segments as functionally different safety domains when modeling near-miss dynamics and designing countermeasures.